FINAL Program for the 3rd International IMACS Conference on Applications of Computer Algebra

July 24, 1997 (Thursday)

8:00 am  Lokelani: ACA’97 Welcome

SESSION Computer Algebra Methods in Integration and Analysis of ODEs: Lie Symmetry Methods Vladimir P. Gerdt

Lokelani Moderated FORUM on Computer Algebra Systems Michael Wester

Lealoha Symbolic Computing Aspects of Multi-Agent Systems and Autonomous Agents John Campbell

Lehua

8:15 am An Algorithm for Constructing Integrating Factors and Corresponding First Integrals for Systems of ODE’s Stephen C. Anco & George W. Bluman* Session starts after the first break … Integrating Computer Algebra Systems

and Theorem Provers

Jacques Calmet

8:45 am Using Lie Groups to Solve Differential Equations Stephen V. Coggeshall

9:15 am  BREAK

9:45 am Three Methods for Exact Integration of ODEs Compared Alexei Bocharov Panel Richard Fateman, Hoon Hong, Robert Lewis, Michael Monagan, Richard Petti, Bill Pletsch, Stanly Steinberg, Michael Wester

10:15 am Moving Coframes

Peter Olver

10:45 am Quasi-exactly Solvable Spectral Problems for Ordinary Differential Operators Niky Kamran

11:15 am Equivalence Classes and Symmetries of Abel’s Equation Fritz Schwarz

11:45 am  LUNCH

SESSION Algorithmic Analysis of PDEs using Elimination Algorithms Greg Reid

Lokelani Non-standard Applications Eugenio Roanes-Lozano

Lealoha Automatic Differentiation Stephen Watt & Christele Faure

Lehua

1:45 pm Introduction to Geometric and Algebraic Algorithms for Nonlinear PDE Gregory Reid Applications of Computer Algebra to Image Understanding, Object Recognition, and Computer Vision Peter F. Stiller Automatic Differentiation in Computer Algebra Systems Michael Monagan

2:15 pm Moving Frames and Calculation of Lie Symmetry of PDE Ian Lisle A Polynomial Model for Multi-Valued Logics with a Touch of Algebraic Geometry and Computer Algebra Eugenio Roanes-Lozano*, Luis M. Laita & E. Roanes-Macias Data Assimilation in Meteorology: A Quest for Automatic Differentiation Francois-Xavier Le Dimet

2:45 pm Applying the Factorization Technique for Simplification of Expressions Involving Differential Operator Inverses Serguei P. Tsarev Progress in Natural Language Query in Macsyma Richard J. Petti

3:15 pm  BREAK

3:30 pm Application of Involutive Division Methods to Polynomial and Differential Equations Vladimir P. Gerdt Nonequilibrium Molecular Dynamics Integration Using Maple Debra J. Searles, Dennis J. Isbister* & Denis J. Evans Automatic Differentiation of Code Christele Faure

4:00 pm Rankings of Partial Derivatives and Elimination Algorithms for PDE Colin Rust Analyzing the Thermodynamics of Steam Reforming with Maple Johannes Grotendorst* & Jürgen Dornseiffer Advanced Algorithms for Derivative Propagation Jason Abate

4:30 pm A Direct Construction Method for Conservation Laws of Partial Differential Equations Stephen C. Anco & George W. Bluman* Social Structures and Speeding Trucks: A Simulation with Experimentation Michael A. Faia To Be Announced Stephen Watt

5:00 pm  Lokelani: ACA Business Meeting

July 25, 1997 (Friday)

SESSION Computer Algebra Methods in Integration and Analysis of ODEs: Linear Equations and Hamiltonian Systems Fritz Schwarz

Lokelani Grand Challenges in Computer Algebra Laurent Bernardin

Lealoha

8:15 am Characteristic Classes for Singularities of Linear Differential Equations A. Levelt The Six-Line Problem and Polynomial Resultants Robert H. Lewis

8:45 am Solving Linear Differential Equations with Polynomial Coefficients in Some Power Series Classes Sergei A. Abramov Parallel Symbolic Computation on the Desk-Top Wolfgang Küchlin

9:15 am On the New Face of Kovacic’s Algorithm Manuel Bronstein The Diatomic Anharmonic Oscillator According to Matrix Mechanics Michael Monagan* & John Ogilvie

9:45 am  BREAK

10:15 am Factorization of Linear Differential Operators Mark van Hoeij Fast Multiprecision Evaluation of Series of Rational Numbers Bruno Haible* & Thomas Papanikolaou

10:45 am Searching Local Periodic Families of ODEs in a Normal Form Method Victor F. Edneral Computing on the Net or: Computer Algebra goes Java Laurent Bernardin

11:15 am On Standard Bases over Newton’s Polytopes:

Generalization of Buchberger’s Algorithm

Victor F. Edneral* & Nikolay N. Vassiliev Parallel Polynomial Real Root Isolation Werner Krandick

11:45 am  LUNCH

SESSION Algorithmic Computation of Conservation Laws Greg Reid, Thomas Wolf & Alexei Bocharov

Lokelani Grand Challenges in Computer Algebra Laurent Bernardin

Lealoha Computer Algebra in Education Ian Cohen & Martin Lesser

Lehua

1:45 pm A Comparison of Two Methods to Calculate Conservation Laws of PDE systems Thomas Wolf Factoring Univariate Polynomials over Finite Fields with LiDIA Thomas Pfahler Teaching Basic Mechanics With Computer Algebra, Sophia and Maple Martin Lesser

2:15 pm Symbolic Computation of Conserved Densities for Nonlinear Evolution and Differential-Difference Equations Ünal Göktas* & Willy Hereman Parallel Quantifier Elimination Hoon Hong, Richard Liska, Nicolas Robidoux & Stanly Steinberg Theorema: Application of Mathematica for Proof Training Bruno Buchberger* & Tudor Jebelean

SESSION Software Integration

Olaf Bachmann & Andreas Sorgatz

2:45 pm Symmetric Conservation Laws Mark Hickman The MP protocol: Past, Present, Future Simon Gray Improving Math Teaching with DE-RIVE and the TI-92: The Scaffolding Method Bernhard Kutzler

3:15 pm  BREAK

SESSION Computer Algebra in Robotics

Jochen Pfalzgraf

3:30 pm Some Reflections on Knowledge Based System Modeling - Mathematical Aspects: Applications to Symbolic (and Geometric) Modeling of Robot Arms

Jochen Pfalzgraf Experiences on Software Integration Based on MP Olaf Bachmann Computer Algebra Systems in Beginning College Mathematics Instruction Bill Pletsch

4:00 pm Dynamically Linked Modules of Mu-PAD Andreas Sorgatz

Teaching Discrete Mathematics with Computer Algebra: A New Approach to Modern Algebra and Error-Correcting Codes

Kenneth Hulth* & Igor Gachkov

4:30 pm Using the ‘Ausdehnungslehre’ and ‘Mathematica’ to Determine the Unknown Parameters of Grassmann Chain Mechanisms with Two and Three Input Sources

Gloria Bitterfeld Accomplishing Software Integration: Dynamic Linking vs Interprocess Communications Olaf Bachmann, Simon Gray, Andreas Sorgatz & Hannes Schönemann

FINAL Program for the 3rd International IMACS Conference on Applications of Computer Algebra

July 26, 1997 (Saturday)

SESSION Computer Algebra Methods in Integration and Analysis of ODEs: Approximate Solutions Vladimir P. Gerdt & Fritz Schwarz

Lokelani Problem Solving Environments/Symbolic-Numeric Interfaces Bruce Char, Jeremy Johnson, Sanjiwa Weerawarana &

Lealoha

8:30 am Computer Algebra and the Method of Modified Equations Mohammad O. Ahmed & Robert M. Corless* Problem Solving Environments for Airshed Models Donald Dabdub

9:00 am Nonlinear Differential Equations with Inexactly Known Boundary Conditions and Gröbner bases Vladimir P. Gerdt*, Michael G. Dmitriev & Marina V. Nesterova The PSEware Project: Infrastructure for Building Problem-Solving Environments Randy Bramley

9:30 am Concepts and Algorithms of Computer Analysis for an Approximate Solution of ODE’s Karl Hantzschmann The Use of Algorithm Templates in a Problem Solving Environment for the Numerical Solution of Differential Equations R. L. Akers, Elaine Kant, C. J. Randall, Stanly Steinberg* & R. L. Young

10:00 am  BREAK

10:30 am Point Mapping-Based Analysis of Nonlinear Systems Using Symbolic Algebra Ramesh S. Guttalu* & Henryk Flashner Vertically Averaged Stratified Fluid Flow Models - Analysis and Numerical Solution Supported by Computer Algebra Tools Richard Liska* & Burton Wendroff

11:00 am Curvature in Computer Vision Cheri Shakiban Algorithm Stabilization Techniques and their Application to Symbolic Computation of Generalized Inverses Hiroyuki Minakuchi, Hiroshi Kai, Kiyoshi Shirayanagi & Matu-Tarow Noda*

11:30 am Panel Discussion: The Role of Computer Algebra in PSEs Panel